OpenStudy (anonymous):

Are the lines L1 and L2 passing through the given pairs of points parallel, perpendicular, or neither parallel nor perpendicular? L1: (1,2) , (3,1) and L2: (0,-1) , (2,0)

OpenStudy (tkhunny):

Have you considered calculating the slope of each line?

OpenStudy (anonymous):

idk how to do that

OpenStudy (texaschic101):

we need to find the slopes using slope formula : slope(m) = (y2 - y1) / (x2 - x1) (1,2) x1 = 1 and y1 = 2 (3,1) x2 = 3 and y2 = 1 now we sub slope(m) = (1 - 2) / (3 - 1) m = -1/2 now you find the slope of the other two sets of points

OpenStudy (texaschic101):

let me start you off... slope(m) = (y2 - y1) / (x2 - x1) (0,-1) x1 = 0 and y1 = -1 (2,0) x2 = 2 and y2 = 0 now we sub slope(m) = (0 - (-1) / (2 - 0) can you finish and find the slope ?

OpenStudy (texaschic101):

hello ??

OpenStudy (anonymous):

sorry i took long but i got it now :)